OpenStudy (anonymous):

Factor the difference of two cubes x third-729

OpenStudy (saifoo.khan):

how about leaving some for tommorrow?

OpenStudy (saifoo.khan):

\[(x-9) (x^2+9 x+81) \]

OpenStudy (anonymous):

it is due tonight...was in a wreck Friday evening and just got back home today from hosp so did not get to work on it through out weekend

OpenStudy (saifoo.khan):

Ohh.

OpenStudy (anonymous):

perfect cubes: \[A^3+B^3= (A+B)(A^2-AB+B^2)\] \[A^3-B^3=(A-B)(A^2+AB+B^2)\]